OpenStudy (please.help.me):

Find the exact value of the trigonometric expression given that sin u = 5/13 and cos v =− 3/5 (Both u and v are in Quadrant II.) tan(u + v)

OpenStudy (tkhunny):

Knowing that u and v are in Quadrant II, where does that put u + v? It's always good to know before we start looking for things.

OpenStudy (please.help.me):

I understand that part )

OpenStudy (please.help.me):

Tan is negative

OpenStudy (tkhunny):

It is of great value to TELL us where you are so we don't have to guess. Are you sure? \(\pi/2 < u < \pi\) \(\pi/2 < v < \pi\) \(\pi < u +v < 2\pi\) - From this, it could go either way. What's your plan? Calculate the two tangents and use the tangent of a sum formula or use the sine and cosine sum functions and calculate that tangent?

OpenStudy (please.help.me):

Is it \[\frac{ (-5/12) +(4/-3) }{ 1-(-5/12)(4/-3) }\] Is this right so far?

OpenStudy (tkhunny):

So, you are calculating the two individual tangents. How did you do that?

OpenStudy (please.help.me):

I found it through reference triangles and SOH CAH TOA

OpenStudy (please.help.me):

@3mar Can you check my work so far?

OpenStudy (tkhunny):

-5/12 = tan(u) -- That's good.

OpenStudy (tkhunny):

-4/3 = tan(v) -- That's good.

OpenStudy (tkhunny):

Your arithmetic looks fine.

OpenStudy (please.help.me):

I got 63/16 for my final answer and it is wrong! I am so worried now because this homework is due before 12:00. Please tell me where I went wrong

OpenStudy (tkhunny):

I thought you said it was negative?

OpenStudy (please.help.me):

I guess I mistyped! Thank you!!!

OpenStudy (3mar):

Sorry I will be back within 45 min Salam!

OpenStudy (please.help.me):

@3mar I got it now! Thanks though!)

OpenStudy (3mar):

Ok. That is good!